OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

terenzreignz (terenzreignz):

To understand this question, we need the concept of slopes. What is the slope of the line y = 2x + 2 ?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

The slope of the perpendicular line will be the negative reciprocal of the slope in the given equation. Then you just need to write the equation for the perpendicular as y = mx + b, where m is the slope (negative reciprocal of the slope of the given equation), and then plug in the given point for x and y and then solve for b. That will give you the equation of the perpendicular going through the given point. @manueld167

OpenStudy (anonymous):

y = 2x + 6 y = −x + 3 y = −x + 6 y = 2x + 3?

OpenStudy (anonymous):

-1/2x*

OpenStudy (anonymous):

Yes the slope will be -1/2. So our equation of the perpendicular line will look something like y = -1/2x + b. We don't know what be is though, but we know that the line goes through (6, 3). So plug in 6 for x and 3 for y and then solve for b. After you have b, you will have the complete equation.